# Neural Differential Algebraic Equations

### v1

@@ Abstract
Differential-Algebraic Equations (DAEs) describe the temporal evolution of systems that obey both differential and algebraic constraints. Of particular interest are systems that contain implicit relationships between their components, such as conservation relationships. Here, we present Neural Differential-Algebraic Equations (NDAEs) suitable for data-driven modeling of DAEs. This methodology is built upon the concept of the Universal Differential Equation; that is, a model constructed as a system of Neural Ordinary Differential Equations informed by theory from particular science domains. In this work, we show that the proposed NDAEs abstraction is suitable for relevant system-theoretic data-driven modeling tasks. Presented examples include (i) the inverse problem of tank-manifold dynamics and (ii) discrepancy modeling of a network of pumps, tanks, and pipes. Our experiments demonstrate the proposed method’s robustness to noise and extrapolation ability to (i) learn the behaviors of the system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the system.
@@ I INTRODUCTION
Modeling and simulation of systems via Differential-Algebraic Equations (DAEs) can be difficult, both in terms of formulation as well as numerical implementation. Common issues include (i) over- or under-constraining states and (ii) incorrect model specification, either of which can lead to a prohibitively challenging modeling task [1]. While many such algebraic relationships may be known (e.g., mass is conserved through a pipe), the relationships between system states and these algebraic constraints are often difficult to reconcile in a modeling framework (see, for example, index of a DAE and choosing appropriate initial conditions [1]). Incorporating these algebraic relationships into well-established data-driven modeling paradigms is also non-trivial. Here, measurement noise, partial observations, and missing or incorrect physics all confound the ability to effectively build models that obey algebraic constraints.
Furthermore, it is generally rare to construct a DAE-based model of an engineering system from first principles. More common is to adopt simplified models that can capture improperly specified or missing physics at a given fidelity. A common approach is to employ a closure model [2], i.e., a surrogate for the full-fidelity physics inserted into a lower-fidelity model. Closure models have enabled major advancements in scientific computing but still traditionally rely upon rigorous reduction from first principles physics to derive required simplified relationships [3].
To address this issue, we present a framework for the data-driven modeling of DAEs with a novel neural timestepper scheme. Our framework leverages themes from operator splitting methods [4] and Physics-Informed Neural Network (PINN)-style [5] optimization problems to accommodate algebraic relationships. We demonstrate the utility of these methods on two closure modeling tasks: (i) a tank-manifold problem, where one needs to ‘invert’ the model to provide an estimate of a property of a model component, and (ii) a tank-pipe-pump network for learning unknown nonlinear interactions. Our specific contributions are:
@@
  • We demonstrate the utility of the method in applications to canonical closure and inverse modeling of constrained networked system dynamics.
The paper is organized as follows. In Section II, we give a brief overview of recent methods proposed for data-driven modeling of agent-based systems. Section III details our proposed methodology and its implementation. Example results are provided in Section IV, followed by a discussion in Section V.
@@ II-A Differential-Algebraic Equations
$\begin{split}\frac{dx}{dt}&=f\left(x,y,u\right)\\ 0&=g\left(x,y,u\right),\end{split}$ (1)
@@
A standard DAE solution approach is to transform the system into a set of ODEs to be integrated with standard numerical solvers. This process is termed index reduction [1]. The index of a DAE is a distance measure between a DAE and its associated ODE: an index-1 DAE requires one differentiation of algebraic relationships to yield a consistent set of ODEs, index-2 requires two differentiations, etc. After differentiation and substitution into other system relationships, the governing equations can be written as a set of first-order ODEs.
@@ II-B Neural Differential Equations
where the superscript contained in parentheses denotes indexing by time, not exponentiation. Practically, this integration is performed with a standard numerical ODE solver, which we denote as:
@@
A fundamental assumption of a NODE model is that a direct relationship exists between each system state and its temporal derivative. This is a reasonable assumption for many cases, but poses challenges for partially-observed systems [11] and application to DAEs. Index reduction is a reliable technique for recasting a DAE into a set of ODEs for standard numerical integration [12], but this is incompatible with the present data science problem. When governing equations are unknown or incomplete, the equivalent set of ODEs and required state augmentations are also unknown.
II-C Fractional Step Methods
Our methodology is inspired by fractional step methods for evolving complex multiphysics problems. In fractional step methods (also termed ‘Operator Splitting’), the evolution of a system is split into simpler, independent sub-tasks. The underlying ODE or PDE is then solved explicitly as the sequential solution of these sub-tasks, albeit at a decreased level of accuracy (i.e., splitting error [13]). There are several well-established mathematical techniques for decomposing the solution of an ODE into sub-tasks. The simplest scheme, Lie-Trotter operator splitting [4], assumes a differential operator can be decomposed into two sub-tasks. For the ODE of Eq. 2, assume $f$ can be decomposed into operators $A$ and $B$ such that:
$\frac{dx}{dt}=A(x)+B(x)$ (5)
@@ II-C Fractional Step Methods
$\hat{x}^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(A(x),x^{(t)}\right)$ (6)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(B(x),\hat{x}^{(t+\Delta t)}\right)$ (7)
@@
The first sub-task solves the ODE defined by the operator $A(x)$ on the time interval $[t,t+\Delta t]$. The second sub-task solves the ODE defined by the operator $B(x)$ on the same time interval, but with the initial conditions defined by the solution of the first sub-task. Many variations of this scheme exist for different applications and orders of accuracy. For example, Strang splitting [14] is a second-order accurate scheme that uses half-steps (i.e., steps of $\frac{\Delta t}{2}$) in the above scheme.
II-D Related Work
II-D1 Machine Learning for Dynamical Systems
Methods for learning dynamics from data exist on a continuum from ‘black-box’ models to ’white-box’ models [15]. The ‘black-box’ modeling task is performed without regard to the underlying system or domain knowledge and includes Recurrent Neural Networks (RNNs) and their variants (e.g., LSTMs)  [16], neural state space models [17], neural timesteppsers  [8], sparse symbolic regression [18], and Neural ODEs  [9]. Often, such models extract salient features in the training data to infer future states. However, they are at risk of poor generalization (overfitting) and lack of interpretability, except for symbolic regression techniques [18]. ‘White-box’ models use domain knowledge to constrain models to match expected dynamics by construction and often lead to parameter-tuning problems [19]. In ‘gray-box’ models, recently referred to as physics-informed machine learning [20], prior knowledge is incorporated in a manner to constrain a ‘black-box model’. Neural ODEs constructed with domain knowledge, termed Universal Differential Equations  [6, 21], are an attractive modeling platform for these problems because of their flexibility in encoding physics and other structural priors.
II-D2 Data-driven Modeling of DAEs
If the DAE is well-characterized, the parameter estimation, inverse design, and control tasks are possible through an adjoint sensitivity analysis  [22, 23, 24]. This strategy is generally successful if (i) the index of the underlying DAE is low and (ii) all constituent physics are known.
Recent works have begun to address some common issues regarding data-driven DAE-based models. Moya and Lin  [25] use explicit Runge-Kutta time-stepping to combine semi-explicit DAEs into machine learning frameworks, which they demonstrate by modeling power networks. Xiao et al.  [26] construct neural ODE and DAE modules to forecast a regional power network. Huang et al.  [27] adopt a simplified approach to the construction of a surrogate system for sequence-to-sequence mappings for DAEs. These studies share features of expressing the learning task in terms of a semi-explicit DAE with varying degrees of parameterization and integration of domain knowledge. In addition to these studies, Zhong et al.  [28] addressed contact modeling in multi-body dynamics, the authors in the work handled a constraint update architecture with similar explicit time-stepping as in Moya  [25] and Xiao  [26].
While the specific treatment of each of these domain-focused problems varies, each leverages a form of a multi-step integration technique that separates the evolution of differential and algebraic states. Here, we recognize that the commonality of these works holds the potential for a more powerful generalization of standard neural integration. Specifically, we recast these ideas under the umbrella of fractional-step methods, where integration is split into simpler, sequential sub-tasks. In effect, we propose a new class of neural integrators that are well-suited for a wide variety of modeling tasks, including DAEs, boundary value problems, and other multi-physics systems.
III METHODS
@@
Consider the DAE system of Eq. 1. Similar to the approach of a Lie-Trotter fractional step scheme, we split the evolution of a system’s differential and algebraic states into sequential sub-tasks. The first sub-task updates the algebraic states using a trainable deep neural network (DNN). The second sub-task takes as input the updated algebraic states and performs a single-step integration. These two sequential sub-tasks define one complete forward pass; all system states are updated exactly once. More formally, consider a differentiable ODE layer, denoted as $\operatorname{ODESolve}$. Using a fractional step method as our solution technique for Eq. 1, we formulate a Neural DAE (NDAE) as:
$\begin{split}y^{(t+\Delta t)}&=h\left(x^{(t)},y^{(t)},u^{(t)};\theta_{h}\right)\\ x^{(t+\Delta t)}&=\operatorname{ODESolve}\left(f,\{x^{(t)},y^{(t+\Delta t)},u^{(t)}\};\theta_{f}\right),\end{split}$ (8)
@@
Note that in this formulation, the updates to the algebraic and differential states are sequential; that is, the $\operatorname{ODESolve}$ step is dependent upon the update to the algebraic states. The sub-tasks do not commute. This is an essential assumption that enables explicit updates for both algebraic states and differential states. Figure 1 depicts the handling of differential and algebraic states through the Neural DAE. The module generates an internal state $y^{(t+\Delta t)}$ that is processed by the subsequent $\operatorname{ODESolve}$ as an exogenous input, similar to the notion of Neural Controlled Differential Equations  [29].
@@
Figure 1: Schematic of the proposed DAE Timestepper.
Training the algebraic update sub-task is accomplished by formulating a PINN-style penalty loss [5] in which known constraints are used to ‘penalize’ the discrepancy between constraint satisfaction and predicted states. These penalties are assessed in a multi-objective optimization loss function:
@@
where $\mathcal{L}_{\text{residual}}$ is a measure of the residual between the training dataset and the model’s output, $\mathcal{L}_{\text{constraints}}$ is a regularization term penalizing the constraint violations. The $\lambda$’s represent weightings for balancing the multi-objective loss function. We optimize model parameters $\Theta$ with an off-the-shelf stochastic gradient-based optimizer [30] to solve $\Theta^{*}=\underset{\Theta}{\text{argmin}}~{}\mathcal{L}(\Theta)$.
For this work, we choose minimal architectures for ease of implementation and exposition, as outlined in Alg. 1. We use a first-order explicit timestepper for the ODE integration and use a DNN as a neural surrogate of an algebra solver.
@@
Parameters : $\theta_{f},\theta_{h}$: Tunable parameters for mappings $f$ and $h$

### v2

# Learning Neural Differential Algebraic Equations via Operator Splitting

@@ Abstract
Differential-Algebraic Equations (DAEs) describe the temporal evolution of systems that obey both differential and algebraic constraints. Of particular interest are systems that contain implicit relationships between their components, such as conservation relationships. Here, we present an Operator Splitting (OS) numerical integration scheme for learning unknown components of Differential-Algebraic Equations from time-series data. This methodology is built upon the concept of the Universal Differential Equation; that is, a model constructed as a system of Neural Ordinary Differential Equations informed by theory from particular science domains. In this work, we show that the proposed OS-based time-stepping scheme is suitable for relevant system-theoretic data-driven modeling tasks. Presented examples include (i) the inverse problem of tank-manifold dynamics and (ii) discrepancy modeling of a network of pumps, tanks, and pipes. Our experiments demonstrate the proposed method’s robustness to noise and extrapolation ability to (i) learn the behaviors of the system components and their interaction physics and (ii) disambiguate between data trends and mechanistic relationships contained in the system.
@@ I INTRODUCTION
Modeling and simulation of systems via Differential-Algebraic Equations (DAEs) can be difficult, both in terms of formulation as well as numerical implementation. Common issues include (i) over- or under-constraining states and (ii) incorrect model specification, either of which can lead to a prohibitively challenging modeling task [1]. While many such algebraic relationships may be known (e.g., mass is conserved through a pipe), incorporating these algebraic relationships into well-established data-driven modeling paradigms is non-trivial. Here, measurement noise, partial observations, and missing or incorrect physics confound the ability to effectively build data-driven models that obey algebraic constraints. In practice, simplified models can be adopted that capture improperly specified or missing physics at a given fidelity. A common approach is to employ a closure model [2], i.e., a surrogate for the full-fidelity physics inserted into a lower-fidelity model. Closure models have enabled major advancements in scientific computing but still traditionally rely upon rigorous reduction from first principles physics to derive required simplified relationships [3].
I-A Contributions
To address this issue, we present a novel framework for the data-driven modeling of DAEs. Our framework integrates ideas from neural timesteppers [4], operator splitting methods [5], and Physics-Informed Neural Networks (PINNs) [6] to accommodate algebraic relationships in the context of neural ordinary differential equations (NODEs) [7]. We demonstrate the utility of these methods on two data-driven closure modeling tasks: (i) a tank-manifold problem, where one needs to ‘invert’ the model to provide an estimate of a property of a model component, and (ii) a tank-pipe-pump network for learning unknown nonlinear interactions. Our specific contributions are:
@@ I-A Contributions
  • We demonstrate the utility of the method in data-driven parameter estimation problems such as canonical closure and inverse modeling of constrained networked system dynamics.
  • We open-source the implementation of the proposed method  [^1] https://github.com/pnnl/NeuralDAEs to promote reproducibility and adoption.
The paper is organized as follows. Section II provides a brief background. Section III defines the problem of interest. Section IV details our proposed methodology and its implementation. Example results are provided in Section V, followed by a discussion in Section VI.
I-B Related Work
I-B1 Machine Learning for Dynamical Systems
Methods for learning dynamics from data exist on a continuum from ‘black-box’ models to ’white-box’ models [8]. The ‘black-box’ modeling task is performed without regard to the underlying system or domain knowledge and includes Recurrent Neural Networks (RNNs) and their variants (e.g., LSTMs) [9], neural state space models [10, 11], neural timesteppsers [4], sparse symbolic regression [12], and Neural ODEs (NODEs) [7, 13, 14]. Often, such models extract salient features in the training data to infer future states. However, satisfying algebraic constraints in these data-driven models is not trivial. ‘White-box’ models use domain knowledge to constrain models to match expected dynamics, often leading to parameter-tuning problems [15]. In ‘gray-box’ models, recently referred to as physics-informed machine learning [16], prior knowledge is incorporated in a manner to constrain a ‘black-box model’. Neural ODEs constructed with domain knowledge, termed Universal Differential Equations [17, 18], are an attractive modeling platform for these problems because of their flexibility in encoding structural priors.
I-B2 Data-driven Modeling of DAEs
If the DAE is well-characterized, the parameter estimation, inverse design, and control tasks are possible through an adjoint sensitivity analysis [19, 20, 21]. This strategy is generally successful if (i) the index of the underlying DAE is low and (ii) all constituent physics are known. Recent works have begun to address some common issues regarding data-driven DAE-based models. Moya and Lin [22] use explicit Runge-Kutta time-stepping to combine semi-explicit DAEs into machine learning frameworks, which they demonstrate by modeling power networks. Xiao et al. [23] construct neural ODE and DAE modules to forecast a regional power network. Huang et al. [24] adopt a simplified approach to the construction of a surrogate system for sequence-to-sequence mappings for DAEs. These studies share features of expressing the learning task in terms of a semi-explicit DAE with varying degrees of parameterization and integration of domain knowledge. In addition to these studies, Zhong et al. [25] addressed contact modeling in multi-body dynamics, the authors in the work handled a constraint update architecture with similar explicit time-stepping as in Moya [22] and Xiao [23].
While the specific treatment of each of these domain-focused problems varies, each leverages a form of a multi-step integration technique that separates the evolution of differential and algebraic states. We recognize that the commonality of these works holds the potential for a generalization of standard NODE integration. Specifically, we recast these ideas under operator splitting methods, where integration is split into simpler, sequential sub-tasks. In effect, we propose a new class of neural timesteppers suitable for the data-driven modeling of DAEs with partially unknown dynamics.
@@ II-A Differential-Algebraic Equations
$\begin{split}\frac{dx}{dt}&=f\left(x,y,u\right),\\ 0&=g\left(x,y,u\right),\end{split}$ (1)
@@
A standard DAE solution approach is to transform the system into a set of ODEs to be integrated with standard numerical solvers. This process is termed index reduction [1]. The index of a DAE is a distance measure between a DAE and its associated ODE: an index-1 DAE requires one differentiation of algebraic relationships to yield a consistent set of ODEs, an index-2 DAE requires two differentiations, etc. After the index reduction, the governing equations can be written as a set of first-order ODEs.
@@ II-B Neural Differential Equations
where the superscript contained in parentheses denotes indexing by time. This integration is performed with a standard numerical ODE solver, which we denote as:
@@
III PROBLEM FORMULATION
A fundamental assumption of a NODE model (2) is that a direct relationship exists between each system state and its temporal derivative. This is a reasonable assumption for many cases, but poses challenges for partially-observed systems [28] and application to DAEs. Index reduction is a reliable technique for recasting a DAE into a set of ODEs for standard numerical integration [29], but this is incompatible with the problem when the DAE governing equations are unknown or incomplete, leading to an unknown system of ODEs and required state augmentations.
To address this limitation, we aim to solve the following parameter estimation problem for neural DAEs (NDAE):
$\underset{\Theta}{\text{min}}$ $\int_{t_{0}}^{t_{N}}\Big{(}||x-\hat{x}||_{2}^{2}+||y-\hat{y}||_{2}^{2}\Big{)}$ (5a)
s.t $\frac{dx}{dt}=f\left(x,y,u;\Theta\right),$ (5b)
$0=g\left(x,y,u;\Theta\right),$ (5c)
$x({t_{0}})=x_{0},\,\,\,y({t_{0}})=y_{0},$ (5d)
Where $\hat{x},\hat{y}$ represent reference state variables generated by the unknown DAE system (1). The objective of the problem (5a) is to train the parameters $\Theta$ such that the reference state trajectories in the time interval $[{t_{0}},{t_{N}}]$ are closely approximated by the NDAE model given via Eq. (5b) and Eq. (5d), respectively.
Assumptions 1
Lets consider the unknown DAE system (1) with $f$ and $g$ being continuously differentiable functions in a neighborhood of the point $(x_{0},y_{0},u_{0})$. Suppose the following conditions hold:
  1. 1.
    (Regularity) The Jacobian matrix $\frac{\partial g}{\partial y}$ is nonsingular at $(x_{0},y_{0},u_{0})$, i.e., $\det\big{(}\frac{\partial g}{\partial y}(x_{0},y_{0},u_{0})\big{)}\neq 0$.
  2. 2.
    (Consistency) The initial values satisfy the algebraic constraint, i.e., $g(x_{0},y_{0},u_{0})=0$.
Theorem III.1 (Picard-Lindelöf for index-1 DAEs [30])
Under the above assumptions, there exists a unique, continuously differentiable solution $(x^{(t)},y^{(t)})$ defined on an interval $[t_{0},t_{0}+\Delta t)$, for some $\Delta t>0$, such that the DAE system is satisfied and $(x^{(t_{0})},y^{(t_{0})})=(x_{0},y_{0})$. Furthermore, the algebraic constraint can be locally solved for $y$ as a function of $x$ and $u$, thereby reducing DAEs into a system of ODEs $\frac{\partial x}{\partial t}=f(x,h(x,u),u)$ to which the classical Picard-Lindelöf Theorem III.2 applies.
Theorem III.2 (Classical Picard-Lindelöf for ODEs [31])
Let $f:D\subseteq\mathbb{R}^{n_{x}}\times\mathbb{R}^{n_{u}}\to\mathbb{R}^{n_{x}}$ be a function that is continuous in both arguments and Lipschitz continuous with respect to the state $x$, uniformly in $t$, on an open set $D$ containing the point $(t_{0},x_{0})$. Then there exists an interval $I=[t_{0}-\epsilon,t_{0}+\epsilon]\subset\mathbb{R}$, for some $\epsilon>0$, and a unique continuously differentiable function $:I\to\mathbb{R}^{n_{x}}$ such that
$\frac{\partial x(t)}{\partial t}=f(x(t),u(t)),\quad x(t_{0})=y_{0},$
for all $t\in I$. In other words, the ODE initial value problem has a unique local solution.
IV METHOD
In this paper, we propose an efficient operator splitting (OS) numerical integration scheme for the neural DAE parameter estimation problem given by Eq. (5a). Specifically, we utilize the above Picard-Lindelöf Theorem III.1 to derive a novel operator splitting-based numerical integration method for data-driven modeling of index-1 DAEs.
IV-A Operator Splitting Methods
Our methodology is inspired by operator splitting methods (OS) for evolving complex multiphysics problems. In OS, also called fractional step methods, the evolution of a system is split into simpler, independent sub-tasks. The underlying ODE is then solved explicitly as the sequential solution of these sub-tasks, albeit at a decreased level of accuracy (i.e., splitting error [32]). There are several well-established techniques for decomposing the solution of an ODE into sub-tasks. The simplest scheme, Lie-Trotter operator splitting [5], assumes a differential operator can be decomposed into two sub-tasks. For the ODE of Eq. (2), assume $f$ can be decomposed into operators $A$ and $B$ such that:
$\frac{dx}{dt}=A(x)+B(x),$ (6)
@@ IV-A Operator Splitting Methods
$\hat{x}^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(A(x),x^{(t)}\right),$ (7)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(B(x),\hat{x}^{(t+\Delta t)}\right).$ (8)
@@
The first sub-task solves the ODE defined by the operator $A(x)$ on the time interval $[t,t+\Delta t]$. The second sub-task solves the ODE defined by the operator $B(x)$ on the same time interval, but with the initial conditions defined by the solution of the first sub-task. Many variations of this scheme exist for different applications and orders of accuracy. For example, Strang splitting [33] is a second-order scheme that uses half-steps (i.e., steps of $\frac{\Delta t}{2}$) in the above scheme.
IV-B Operator Splitting for Differential Algebraic Equations
@@
Now consider the DAE system of Eq. (1). Similar to the approach of a Lie-Trotter splitting scheme, we split the evolution of a system’s differential and algebraic states into sequential sub-tasks. The first sub-task updates the algebraic states using a trainable deep neural network (DNN). The second sub-task takes as input the updated algebraic states and performs a single-step integration. These two sequential sub-tasks define one complete forward pass; all system states are updated exactly once. More formally, consider a differentiable ODE layer, denoted as $\operatorname{ODESolve}$. Using the operator splitting method as our solution technique for Eq. (1), we formulate a time stepper for Neural DAE as:
$y^{(t+\Delta t)}$ $=h\left(x^{(t)},y^{(t)},u^{(t)};\theta_{h}\right)$ (9a)
$x^{(t+\Delta t)}$ $=\operatorname{ODESolve}\left(f,\{x^{(t)},y^{(t+\Delta t)},u^{(t)}\};\theta_{f}\right),$ (9b)
@@
Note that in this formulation, the updates to the algebraic and differential states are sequential; that is, the $\operatorname{ODESolve}$ step is dependent upon the update to the algebraic states. The sub-tasks do not commute. This is an essential assumption that enables explicit updates for both algebraic states and differential states. Figure 1 depicts the handling of differential and algebraic states through the proposed OS-based NDAE integration scheme. The module generates an internal state $y^{(t+\Delta t)}$ that is processed by the subsequent $\operatorname{ODESolve}$ as an exogenous input, similar to the notion of Neural Controlled Differential Equations [34].
@@
Figure 1: Schematic of the proposed operator splitting-based DAE Timestepper.
IV-C Loss Function and Optimization Algorithm
In this paper, solving the algebraic update sub-task is accomplished by formulating a PINN-style penalty loss [6] in which known constraints are used to ‘penalize’ the discrepancy between constraint satisfaction and predicted states. These penalties are assessed in a multi-term loss function:
@@ IV-C Loss Function and Optimization Algorithm
where $\mathcal{L}_{\text{residual}}(X,\hat{X};\Theta)=\sum_{t=0}^{N}\left(||x_{t}-\hat{x}_{t}||_{2}^{2}+||y_{t}-\hat{y}_{t}||_{2}^{2}\right)$ is a measure of the residual between the model’s differential and algebraic states ${X}=\{({x}^{(0)},{y}^{(0)}),...,({x}^{(N)},{y}^{(N)})\}$, and the training dataset $\hat{X}=\{(\hat{x}^{(0)},\hat{y}^{(0)}),...,(\hat{x}^{(N)},\hat{y}^{(N)})\}$, where $N$ defines the length of a finite-horizon rollout. The second term $\mathcal{L}_{\text{constraints}}(X,\hat{X};\Theta)=\sum_{t=0}^{N}||g(x_{t},y_{t},u_{t})||^{2}_{2}$ is penalizing the constraint violations. The $\lambda$’s represent weightings for balancing the multi-objective loss function.
We optimize model parameters $\Theta$ with an off-the-shelf stochastic gradient-based optimizer [35] to solve $\Theta^{*}=\underset{\Theta}{\text{argmin}}~{}\mathcal{L}(\Theta)$. For this work, we choose minimal architectures for ease of implementation and exposition, as outlined in Alg. 1. We use a first-order explicit timestepper for the ODE integration and use a DNN as a neural surrogate of an algebra solver. However, the proposed OS-based NDAE scheme is more generic than that and allows the combination of higher-order ODE solvers with iterative algebraic solvers such as Newton’s method.
@@
Parameters : $\Theta=\{\theta_{f},\theta_{h}\}$: trainable parameters for mappings $f$ and $h$;
gradient step-size hyperparameter $\eta$;
loss function error tolerance $\epsilon$